Write an equation of the line satisfying the given conditions.​
Perpendicular to x−5y=4​; ​y-intercept ​(0,−5​)

Found 2 solutions by josgarithmetic, mananth:
Answer by josgarithmetic(39629)

y=mx+b
;
m neede is m=-5----Do you know how?
;
y=-5x-5

Answer by mananth(16946)

1 x + -5 y = 4
Find the slope of this line

-5 y = -1 x + 4
Divide by -5
y = 1/ 5 x + - 4/5
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1/5

The slope of a line perpendicular to the above line will be the negative reciprocal -5
Because m1*m2 =-1
The slope of the required line will be -5

m= -5 ,point ( 0 , -5 )
Find b by plugging the values of m & the point in
y=mx+b
-5 = -0 + b
b= -5
m= -5
The required equation is y = -5 x -5
m.ananth@hotmail.ca